# XpertAI: uncovering model strategies for sub-manifolds

### v1

@@ 4.2 What sets apart a good wine from an excellent one?
As mentioned in the introduction, we now turn to a slightly more hedonistic and tangible example where many people have strong opinions: the quality of red wine. We utilize Kaggle’s Spanish red wine dataset [^1] https://www.kaggle.com/datasets/fedesoriano/spanish-wine-quality-dataset which contains several thousand wine samples. They are described by five numerical (year, price, as well as body, acidity, and quality scores) and four categorical (name of the winery and the wine, grape, region) features. As a regression target, we have selected the quality score which is an ’average rating’ given by thousands of testers (rating binned into 8 discrete quality levels), and we utilize the remaining eight features as inputs. After data-pre-processing around 1700 samples are left. We have trained a small fully-connected ANN which achieved an $R^{2}$ of around 0.7.
We now want to learn what, according to the model, sets apart a good wine from an excellent one. We define wine as good when it belongs to the top 10 % and excellent when it belongs to the top 1.5 % of the model output range. We train three range experts ($\tau=0.33$) and compare the respective attributions obtained from standard IG with its application within the XpertAI framework. Figure 5 shows the decomposition of the excellent wine attributions into the respective expert bases. Aside from the natural change in attribution scales, the most prominent difference in the explanations is the contribution of the price to the model outcome. For the naive IG attributions, the price is the by far most important feature (meaning high prices alone are the main indicator for excellent wines). The contextualized XpertAI attributions, on the other hand, give a much more balanced picture. Here, the outcome suggests that the price is the most important feature only for the low-quality range expert (meaning what distinguishes an average from a poor wine, blue). The relative importance of the price, however, is significantly reduced when compared to average wines (orange) and almost vanishes when compared to good wines (greed). There, the sum of all other quality criteria is much more important than the price of the wine itself. This directly translates to some actionable (and intuitive) insight: if you next time buy a wine in the supermarket, don’t go cheap to ensure you buy a decent wine. When looking for an excellent one though, you might be better off with the expert judgement of your local wine seller.
Figure 5: Decomposition of naive explanations (grey) for samples from the high output range (’excellent’ wines) with respect to low, medium and high-quality reference values (colorful). The XpertAI explanations allow for nuanced insights into what makes an excellent wine better than the worst (blue), a decent (orange) or a good (green) alternative.
To make sure, our insights are not based on some unfaithful attributions, we also compare quantitative faithfulness for the and observe an average increase in the ABC metric by more than 10 % (see Sec. 5).
@@ 4.3 Why does the wind turbine produce less than expected?
Wind power is one of the pillars of decarbonizing energy systems around the world. Wind turbines are often placed in remote locations and need to be operated and monitored from a distance, using data from their Supervisory Control and Data Acquisition (SCADA) system. Effectively leveraging this data is an active area of research (Innes Murdo Black & Kolios, 2021), with the primary focus on detecting and diagnosing underperformance as the central challenge (Pandit et al., 2023). However, the detection of underperformance is always context-specific since the implicit question is: ‘underperformance relative to what?’ In the wind turbine case, it is the condition without the presence of a malfunction, given the context of prevailing ambient conditions.
We utilize data from a 2 MW wind turbine and a meteorological met-mast from an onshore wind farm on the Iberian peninsula  [^2] https://opendata.edp.com. SCADA data is available for two years and includes ambient conditions as well as technical turbine parameters as 10-minute averaged values (50,000 data points after pre-processing). We have trained a small fully-connected MLP to predict the turbine output from wind speed, air density, and turbulence intensity. The model achieves a competitive RMSE of less than 36 kW. Additionally, we have augmented the data with so-called yaw-misalignment losses. They occur when a turbine does not perfectly face the incoming wind direction, which reduces the effective area of the rotor. Detecting yaw-misalignment is part of ongoing research (Pandit et al., 2021; QU2, 2022) and attributing it by XAI methods has recently been proposed as an effective solution (Letzgus & Müller, 2024). For such an approach to work, we need our XAI methods to faithfully attribute the losses induced by yaw-misalignment to the respective feature (difference of nacelle and wind direction). In our setup, we can directly compare attributions with the respective ground-truth-losses.
@@
We trained five range experts for the different operational regions of the turbine. Figure 6 shows the comparison of attributing the yaw-misalignment induced losses to the respective yaw-feature with standard LRP and our XpertAI-LRP variant. We can observe that the naive LRP application attributes exhibit a systematic overestimation (larger negative values) of losses caused by incorporating phenomena from outside the respective operational regime. Our proposed novel attributions obtained from the range-experts, on the other hand, are on average much closer to the ground truth. For turbine operators, this directly translates to better operation and maintenance decisions and therefore highlights the benefit of using sub-manifold-specific explanations in industrial or engineering applications.
@@ 5 Quantitative Evaluation and Sanity Checks of XpertAI-faithfulness
After having shown some intriguing insights that were enabled by our XpertAI approach in the previous chapter, we now conduct a systematic evaluation of their faithfulness. We utilize the controlled environment of the previously introduced rMNIST dataset to investigate practically relevant questions of how to set up, train, and utilize our range experts. Furthermore, we add some sanity-check type evaluation to empirically validate Proposition 1 from Sec. 3 and get practically relevant insights for range-expert model selection. Details on the respective experiments can be found in Appendix C and some guidance on how to obtain faithful range experts in Appendix B.
@@ 5.1 Are XpertAI attributions faithful?
To answer this question properly, we first need to define how we measure faithfulness. In general, regression-specific evaluation of XAI attribution methods has not received much attention so far, (Hama et al., 2023) being a noteworthy exception. The authors propose a regression-specific flavor of the area under the curve in pixel-flipping (Bach et al., 2015; Samek et al., 2017), called the area between the curves (ABCs). The ABC is defined as the area between the model output when occluding a sample’s features in the order of attribution magnitudes and a straight line connecting $f(x)$ and $f(x^{\prime})$ (which corresponds to random sorting). Since sorting ascending and descending can result in asymmetrical curves, we sum over both areas (Blücher et al., 2024). We furthermore normalize by the distance between the sample and the baseline when averaging. Higher values of ABC are better.
Furthermore, the challenge of including context in attribution methods (Sec. 2.2) naturally extends to occlusion-based evaluation (ergo, what to occlude with?). Here again, the best way to ensure that we evaluate attributions within the relevant output range of function $f(x)$, where we account for context-specific (local) effects, is to occlude with a domain-specific counterfactual (Albini et al., 2022; Hama et al., 2023). Therefore, we sample conditional $x^{\prime}=D(x|y=\widetilde{y})$ from the available data set $D$ with which we then occlude and average the respective ABCs over multiple draws.
Table 1: Comparison of faithfulness for different attribution methods applied naively and within the XpertAI framework. Relative improvement of ABC over naive application.
@@
rMNIST +42 % +7.2 %
WINE +23.6 % +10.6 %
FRIEDMAN +12.7 % +1.9 %
CALIFORNIA +1.8 % +9.7 %
DIABETES +2.6 % +4.4 %
Table 1 reports the ABC scores of our XpertAI approach relative to a naive application of LRP and IG on the previously introduced as well as several popular regression benchmarks (Pedregosa et al., 2011). For each of them, we trained three range experts and evaluated samples from the top range (see Appendix C). Overall, we see consistent improvements in ABC scores across all settings which means that our approach indeed can generate more faithful attributions with respect to a user-specific query. Note, that the advantage is significantly larger for LRP. There, the approach corresponds to a data-driven root-search strategy whereas naively, there is no such option. For IG we have already leveraged its inherent contextualization capability to some extent by utilizing the mean over all input samples as a starting point for the integration path. Our approach is still able to further refine the attributions towards a better contextualization.
5.2 (Diss-)aggregate XpertAI-attributions
From Proposition 1, we can in principle derive an alternative way to obtain disentangled and contextualized attributions with respect to $\widetilde{y}$. Instead of adding up the respective expert attributions, we subtract them from the original explanation in reverse order. Intuitively, only information relevant for higher-range bins should remain. We test this hypothesis empirically on the rMNIST dataset. We flip pixels to zero according to the order of the difference of attributions $\mathcal{E}(y,x)-\sum_{m}{\mathcal{E}(z_{m},x)}$. Intuitively, the more evidence associated with lower-range concepts we subtract, the more evidence for higher values should remain, and therefore the flipping curve should decrease more slowly. In the ideal case, the only information with positive attributions is the one relevant for values larger than $\widetilde{y}_{i}$ and the flipping curve should therefore remain around that value for as long as possible. When testing this empirically, we indeed see such a behavior (Figure 7). Note also, that the plateaus of the different range experts do not cluster around the digit-transitions (0.33 and 0.66). This means, that despite the strong global concept shifts present in the data, the range experts were able to capture more subtle, local effects that guide $f(x)$ in the context of the respective reference values.
Table 2: Results for pixel flipping experiments for regression MNIST. Results within ranges: sample-flipping baseline pairs are within one expert range. ABC values are normalized by flipping distance, high values are better.
@@
5.3 How many range experts?
One practically very relevant question is, how many range experts to train, which includes the choice of their respective ranges ($\tau$). Conceptually, the method should work best if every distinct sub-region of the output is covered by at least one range expert. In practice, these can either be domain-informed and therefore known apriori, or inferred by analyzing activation patterns (from an activation vs. $f(x)$ scatter plot, for example). In the context of our rMNIST case, selecting one range expert for each digit range, therefore three range experts in total appears to be the most intuitive choice. In practice, however, we might not know where exactly these boundaries lay. Therefore, we compare settings with three, five, six, and nine equally spread range experts.
@@ 6 Discussion and Conclusions
In this paper, we have proposed the XpertAI framework to achieve contextualized and disentangled attributions when explaining regression models. Inspired by the MoE approach, the framework divides the data into sub-manifolds, each of which corresponding to a certain predicted output range. Such a division is achieved by building a collection of range experts, which we equip with explainability. It enables for the first time a disentanglement along the output of the prediction strategy and the resolution of specific user-defined queries.
@@
Empirically, we find that our XpertAI framework is able to distill locally relevant explanations from highly aggregated global standard attributions, as demonstrated by a number of quantitative experiments based on occlusion tests. Explanations associated with each expert range can be precomputed, in a way that exact user queries can be answered very quickly as a linear combination of the precomputed explanations.
@@
Our approach can be interpreted within the framework of virtual layers, which has been instrumental in achieving various forms of explanation disentanglement. Furthermore, our approach provides an alternative to the more common approach of extracting reference points or counterfactuals and bypasses some of the challenges associated with reference points, such as their multiplicity and the need to search for them. Also, our approach differs from self-interpretable generalized additive models, by remaining applicable to a broad range of ML models, including deep neural networks.
@@
While we have demonstrated that our method can work alongside various explanation techniques, in particular, gradient-based techniques such as Integrated Gradients, or propagation-based techniques such as LRP, it is necessary for the latter to structurally disentangle the range experts. While we have proposed for this later step a surrogate modeling approach, these surrogates need to be carefully trained and regularized to maintain the original model’s prediction output as well as its prediction strategy. Hybrid approaches, with the top layers handled by perturbation-based techniques and the lower layers with propagation, may eliminate the need for structural disentanglement while at the same time retaining high accuracy and computational efficiency.
@@
Overall, our work has highlighted the need to precisely formulate “what to explain” (the explanandum) and proposed a practical and flexible solution in the context of regression. The MoE idea our method relies on, however, is more general, and our framework could be extended in the future to other decomposition of the predicted output, e.g. for structured output tasks such as time series prediction.
Impact statement: This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Propositions 1 and 2
@@
Appendix B How to train and select good range experts?
In practice, we need to train and select appropriate range experts for the XpertAI approach to enhance contextualization as reported above. This process may vary based on the respective XAI attribution method being employed. For occlusion- and gradient-integration-based methods, which do not require additional structural disentanglement (see Section 3.3), a simple shift-and-clip strategy is sufficient. For propagation-based methods, however, we need to learn the surrogate $a\mapsto(z_{m})_{m}$, as described earlier. Here, we want to highlight the need for appropriate regularization to avoid overfitting, which in the case of range experts would result in unfaithful model attributions. Consequently, we want to choose the least complex range expert, that can sufficiently learn the respective mapping.
@@ Appendix B How to train and select good range experts?
Appendix C Details on Evaluation (Sec. 4 and 5)
Figure 8: Details on datasets, models and range experts for the respective benchmark datasets
Figure 8: Details on datasets, models and range experts for the respective benchmark datasets

### v3

@@ 4.2 What sets apart a good wine from an excellent one?
As noted in the introduction, we now explore a more hedonistic and tangible example - red wine quality. We utilize Kaggle’s Spanish red wine dataset [^2] https://www.kaggle.com/datasets/fedesoriano/spanish-wine-quality-dataset which contains several thousand wine samples. They are described by five numerical (year, price, as well as body, acidity, and quality scores) and four categorical (name of the winery and the wine, grape, region) features. The quality score, which is an ’average rating’ given by thousands of testers (rating binned into 8 discrete quality levels), is our regression target. After data-pre-processing around 1700 samples are left. We have trained a small fully-connected ANN which achieved an $R^{2}$ of around 0.7.
We now want to learn what, according to the model, sets apart a good wine from an excellent one. We define wine as good when it belongs to the top 10 % and excellent when it belongs to the top 1.5 % of the model output range. We train three range experts ($\tau=0.33$) and compare the respective attributions obtained from standard IG with its application within the XpertAI framework. Figure 7 shows the decomposition of the excellent wine attributions into the respective expert bases. Aside from the natural change in attribution scales, the most prominent difference in the explanations is the contribution of the price to the model outcome. For the naive IG attributions, the price is the by far most important feature (meaning high prices alone are the main indicator for excellent wines). The contextualized XpertAI attributions, on the other hand, give a much more balanced picture. Here, the outcome suggests that the price is the most important feature only for the low-quality range expert (meaning what distinguishes an average from a poor wine, blue). The relative importance of the price, however, is significantly reduced when compared to average wines (orange) and almost vanishes when compared to good wines (green). There, the sum of all other quality criteria is much more important than the price of the wine itself. This directly translates to some actionable (and intuitive) insight: if you next time buy a wine in the supermarket, don’t go cheap to ensure you buy a decent wine. When looking for an excellent one though, you might be better off with the expert judgement of your local wine seller.
Figure 7: Decomposition of naive explanations (grey) for samples from the high output range (’excellent’ wines) with respect to low, medium and high-quality reference values (colourful). The XpertAI explanations allow for nuanced insights into what makes an excellent wine better than the worst (blue), a decent (orange) or a good (green) alternative.
To make sure, our insights are not based on intuitive but unfaithful attributions, we also compare quantitative faithfulness for the and observe an average increase in the ABC metric by more than 10 % (see Sec. 5).
@@ 4.3 Why does the wind turbine produce less than expected?
Wind power is one of the pillars of decarbonizing energy systems around the world. Wind turbines are often placed in remote locations and need to be operated and monitored from a distance, using data from their Supervisory Control and Data Acquisition (SCADA) system. Effectively leveraging this data is an active area of research [18], with the primary focus on detecting and diagnosing underperformance as the central challenge [33]. However, the detection of underperformance is always context-specific since the implicit question is: ‘underperformance relative to what operational state?’ In the wind turbine case, it is the condition without the presence of a malfunction, given the context of prevailing ambient conditions.
We utilize data from a 2 MW wind turbine and a meteorological met-mast from an onshore wind farm on the Iberian peninsula  [^3] https://opendata.edp.com. SCADA data is available for two years and includes ambient conditions as well as technical turbine parameters as 10-minute averaged values (50,000 data points after pre-processing). We have trained a small fully-connected MLP to predict the turbine output from wind speed, air density, and turbulence intensity. The model achieves a competitive RMSE of less than 36 kW. Additionally, we have augmented the data with so-called yaw-misalignment losses. They occur when a turbine does not perfectly face the incoming wind direction, which reduces the effective area of the rotor. Detecting yaw-misalignment is an ongoing field of research [34, 37] and attributing it by XAI methods has recently been proposed as an effective solution [25]. For such an approach to work, we need our XAI methods to faithfully attribute the losses induced by yaw-misalignment to the respective feature (difference of nacelle and wind direction). In our setup, we can directly compare attributions with the respective ground-truth-losses.
@@
We trained three range experts across the different operational regions of the turbine (see Appendix). Figure 8 shows the comparison of attributing the yaw-misalignment induced losses to the respective yaw-feature with standard LRP and our XpertAI-LRP variant. We can observe that the naive LRP application attributes exhibit a systematic overestimation (larger negative values) of losses caused by incorporating phenomena from outside the respective operational regime. Our proposed novel attributions obtained from the range-experts, on the other hand, are on average much closer to the ground truth. For turbine operators, this directly translates to better operation and maintenance decisions and therefore highlights the benefit of using sub-manifold-specific explanations in industrial or engineering applications.
@@ 5 Quantitative Evaluation and Sanity Checks of XpertAI-faithfulness
After having presented some intriguing insights enabled through our XpertAI approach in the previous chapter, we now conduct a systematic evaluation of explanation faithfulness. Details on the respective experiments and additional insights for obtaining faithful range experts can be found in Appendix 0.B and 0.C.
@@ 5.1 Are XpertAI attributions faithful?
To answer this question quantitatively, we utilize the ABC score as introduced in Section 3.4. Table 1 reports the ABC scores of our XpertAI approach relative to a naive application of LRP and IG on the previously introduced data sets as well as several popular regression benchmarks [36]. For each of them, we trained three range experts and evaluated samples from the top range (see Appendix 0.C). Overall, we see consistent improvements in ABC scores across all settings which means that our approach indeed can generate more faithful attributions with respect to a user-specific query. Note, that the advantage is significantly larger for LRP where our approach corresponds to a data-driven root-search strategy whereas naively, there is no such option. For IG we have already leveraged its inherent contextualization capability to some extent by utilizing the mean over all input samples as a starting point for the integration path. Our approach is still able to further refine the attributions towards a better contextualization.
Table 1: Comparison of faithfulness for different attribution methods applied naively and within the XpertAI framework. Relative improvement of ABC over naive application. Standard deviation over 5 different retraining runs for LRP.
@@
rMNIST +50.7 % ±3.5 +7.2 %
WINE +19.8 % ±1.2 +10.6 %
FRIEDMAN +12.6 % ±0.4 +1.9 %
CALIFORNIA +2.5 % ±0.9 +9.7 %
DIABETES +3.8 % ±1.6 +4.4 %
5.2 How many range experts?
One practically relevant question is, how many range experts to train, which includes the choice of their respective ranges ($\tau$). Conceptually, the method works best if every distinct sub-region of the output is covered by at least one range expert. In practice, these can either be domain-informed and therefore known apriori, or inferred by analyzing activation patterns (from an activation vs. $f(x)$ scatter plot, for example). In the context of our rMNIST case, selecting one range expert for each digit range, therefore three range experts in total appears to be the most intuitive choice. Since in practice, we might not know where exactly these boundaries lay, we compare settings for three, five, six, and nine equally spread range experts.
Table 2: Results for pixel flipping experiments for regression MNIST. Results within ranges: sample-flipping baseline pairs are within one expert range. ABC values are normalized by flipping distance. Values for naive methods differ because of the normalization. High values are better.
@@ 5.2 How many range experts?
5.3 (Diss-)aggregate XpertAI-attributions
From Proposition 1, we can in principle derive an alternative way to obtain disentangled and contextualized attributions with respect to $\widetilde{y}$. Instead of adding up the respective expert attributions, we subtract them from the original explanation in reverse order. Intuitively, only information relevant to higher-range bins should remain. We test this hypothesis empirically on the rMNIST dataset. We flip pixels to zero according to the order of the difference of attributions $\mathcal{E}(y,x)-\sum_{m}{\mathcal{E}(z_{m},x)}$. Intuitively, the more evidence associated with lower-range concepts we subtract, the more evidence for higher values should remain, and therefore the flipping curve should decrease more slowly. In the ideal case, the only information with positive attributions is the one relevant for values larger than $\widetilde{y}_{i}$ and the flipping curve should therefore remain around that value for as long as possible. When testing this empirically, we indeed see such a behaviour (Figure 9). Note also, that the plateaus of the different range experts do not cluster around the digit-transitions (0.33 and 0.66). This means, that despite the strong global concept shifts present in the data, the range experts were able to capture more subtle, local effects that guide $f(x)$ in the context of the respective reference values.
@@ 6 Discussion and Conclusions
In this paper, we have proposed the XpertAI framework to achieve contextualized and disentangled attributions when explaining regression models. Inspired by the MoE approach, the framework divides the data into sub-manifolds, each of which corresponds to a certain predicted output range. Such a division is achieved by building a collection of range experts, which we equip with explainability. It enables for the first time a disentanglement along the output of the prediction strategy and the resolution of specific user-defined queries.
@@
Empirically, we find that our XpertAI framework can distill locally relevant explanations from highly aggregated global standard attributions, as demonstrated by several quantitative experiments based on occlusion tests. Explanations associated with each expert range can be precomputed, so that exact user queries can be answered very quickly as a linear combination of the precomputed explanations.
@@
Our approach can be interpreted within the framework of virtual layers, which has been instrumental in achieving various forms of explanation disentanglement. Furthermore, our approach provides an alternative to the more common approach of extracting reference points or counterfactuals and bypasses some of the challenges , such as their multiplicity and the need to search for them. Also, our approach differs from self-interpretable generalized additive models, by remaining applicable to a broad range of ML models, including deep neural networks.
@@
We have demonstrated that our method can work alongside various explanation techniques, in particular, gradient-based techniques such as Integrated Gradients, or propagation-based techniques such as LRP. While this enables a seamless integration into existing explanation pipelines our approach naturally inherits potential shortcomings of these methods. Furthermore, it is necessary for propagation-based techniques to structurally disentangle the range experts. While we have proposed a surrogate modeling approach for this step, these surrogates need to be carefully trained and regularized to maintain the original model’s prediction output as well as its prediction strategy. Also, retraining implies additional computational cost. Hybrid approaches, with the top layers handled by perturbation-based techniques and the lower layers with propagation, may eliminate the need for structural disentanglement while at the same time retaining high accuracy and computational efficiency. Enhanced approaches, inspired by model distillation or formally equivalent neural networks, could also be considered.
@@
Overall, our work has highlighted the need to precisely formulate “what to explain” (the explanandum) and proposed a practical and flexible solution in the context of regression. The MoE idea our method builds upon, however, is more general, and our framework could be extended in the future to other decomposition of the predicted output, e.g. for structured output tasks such as time series prediction. Additional future work could furthermore focus on automating the optimal number of experts in a data-driven way. While we have shown that for sufficiently populated ranges of the output adding more experts improves contextualization, there certainly are limitations arising from data availability and computational constraints. Lastly, the application and evaluation to more complex models, such as regression foundation models [16], should be considered in the future.
6.0.1 Acknowledgements
This work was partly funded by the German Ministry for Education and Research [01IS14013A-E, 01GQ1115, 01GQ0850, 01IS18056A, 01IS18025A, and 01IS18037A], the German Research Foundation as Math+: Berlin Mathematics Research Center [EXC2046/1, project-ID: 390685689], the Investitionsbank Berlin [10174498 ProFIT program], and the European Union’s Horizon 2020 Research and Innovation program under grant [965221]. Furthermore, Klaus-Robert Müller was partly supported by the Institute of Information and Communications Technology Planning and Evaluation grants funded by the Korean Government [2019-0-00079]. Our gratitude extends to Jonas Lederer, Pattarawat Chormai and Stefan Blücher for their invaluable comments and feedback that have contributed to enhancing the quality of the manuscript.
Appendix 0.A Proof of Propositions 1 and 2
@@
Appendix 0.B How to train and select good range experts?
In practice, we need to select appropriate range experts for the XpertAI approach to enhance contextualization. This process may vary based on the respective XAI attribution method being employed. For occlusion- and gradient-integration-based methods, which do not require additional structural disentanglement (see Section 3.3), a simple shift-and-clip strategy is sufficient. For propagation-based methods, however, we need to learn the surrogate $a\mapsto(z_{m})_{m}$ (see Sec. 3). Here, we want to highlight the need for appropriate regularization to avoid overfitting, which in the case of range experts would result in unfaithful model attributions. Analogously to regular model selection, we aim to choose the least complex range expert, that can sufficiently learn the respective mapping.
@@ Appendix 0.B How to train and select good range experts?
Appendix 0.C Details on Evaluation (Sec. 4 and 5)
0.C.1 Details face-age regression example
For this analysis, we have made use of a dataset containing $\sim$ 20k facial images associated with biological age [^4] https://www.kaggle.com/frabbisw/facial-age (biased toward younger ages). Each image is pre-processed so that all of them have the same size (200x200) and the faces are aligned and centred. We used a VGG-16 [42] model pre-trained on ImageNet [10, 39] as a feature extractor followed by one ReLU layer with 256 neurons, a dropout-layer, and a final linear layer mapping the 256 neurons to a real-valued age prediction. In all cases we used LRP-$\alpha_{1}\beta_{0}$ rule [6, 32] in the convolutional layers and LRP-$\epsilon$ rule [6] (where biases are ignored) for the fully connected layers.
0.C.2 Details quantitative evaluation
Here, we describe the details of our quanitative experiments. For the rMNIST experiments, we utilized a vanilla CNNs with two convolutional, ReLU and pooling layers, followed by three fully connected layers. The convolutional blocks were kept frozen, and only the fully connected layers were re-trained as experts, starting from their original model weights. The other problems (Wind, Wine, California and Diabetes) are based on tabular data. Here, we utilize a 4 layer-MLP with 20 neurons in each hidden layer. The last two were re-trained for each expert. Moreover, we utilized the PCA initialization trick, described above (Appendix B). $w_{m}$ was selected to be 1 for all expert ranges between reference value and sample output, and 0 otherwise. More specific information on the implementation can be found in the published code repository. [^5] https://github.com/sltzgs/XpertAI